OpenStudy (anonymous):

Find the inverse Laplace transform of e^-5s/(s^2+3s+2)+e^-10s/(s(s^2+3s+2))+(1/2)/(s^2+3s+2). (I've factored the trinomial into (s+1)(s+2). But how do I find the answer? The table doesn't have e^-5s over something as the formula.)

OpenStudy (loser66):

\[\huge\frac{e^{-5s}}{s^2 +3s+2}+\frac{e^{-10s}}{s(s^2+3s+2)}+\frac{\frac{1}{2}}{s^2+3s+2}\] right?

OpenStudy (anonymous):

Right.

OpenStudy (loser66):

I mean \(e^{-5s}\) not \(e^{-5} s\), right? s is in the exponent, right?

OpenStudy (anonymous):

Right.

OpenStudy (loser66):

ok, sweeties, let see whether I can help or not. hihihi

OpenStudy (anonymous):

You want to see the answer?

OpenStudy (loser66):

nope, I need practice too. However, I want original problem to find out f(t)

OpenStudy (anonymous):

What's f(t)?

OpenStudy (loser66):

inverse laplace transform of F(s)

OpenStudy (loser66):

Don't get what I mean? In laplace transform table, you have 2 column, the right one and the left one. The left one is respect to t , that's f(t) and correspond to them is F(s) respect to s. For example, t = 2/s^2 means